# The Pitfalls of Next-Token Prediction

### v2

@@ F.2 Clever Hans Cheating Accuracies
GPT-Mini $99.7$ $100$ $100$ $99.8$ $0.0$
GPT2-Large $99.8$ $99.7$ $100$ $99.8$ $0.0$
Mamba $97.6$ $98.3$ $99.5$ $95.9$ $0.0$
@@ F.3 More Detailed Accuracies
GPT-Mini $49.8$ $49.1$ $19.1$ $8.1$ $0.0$
GPT2-Large $48.9$ $49.2$ $19.4$ $10.3$ $3.5$
Mamba $48.5$ $48.7$ $20.2$ $9.3$ $0.0$
@@
GPT-Mini $99.7$ $99.8$ $100$ $99.8$ $0.0$
GPT2-Large $99.9$ $99.9$ $99.6$ $99.8$ $99.9$
Mamba $98.5$ $96.2$ $99.1$ $99.5$ $0.0$
@@ F.4 Arithmetic Tasks
while keeping all other aspects of the training pipeline (such as the targets and the objective) the same as in standard training. We again train a GPT-Mini-style transformer with AdamW using a learning rate of $5$e-$4$. We display the resulting test accuracies for both standard, reversed and teacherless training in Fig. 6, as a function of training iterations. Here we follow previous works and at every iteration, sample with replacement from all the possible $3$-digit configurations (aside from a test set put aside before). We can indeed see that teacherless training leads to more sample-efficient learning compared to the standard format, but to slightly less efficient learning compared to the reverse format. This again hints at the fact that some form of CleverHans cheating is picked up when learning addition.
In this case though, it is more difficult to precisely characterize the form of the cheat in contrast to the path-finding problem. One possible form of cheating may be that in addition, knowing the leading digits of the answer can provide information about the subsequent digits. For example, when we add two single-digit numbers (picked uniformly at random), knowing that the $10^{\prime}th$ place is a $1$ rather than a 00, tells us that the 00th place is more likely to be a smaller digit like 00.
@@ G.1 Tokenization
We tokenize the graph in the following manner: (1) we first tokenize the randomly shuffled edge list as ``$|v_{1}\hskip 2.84526ptv_{2}|v_{3}\hskip 2.84526ptv_{4}|..."$ where the first vertex in each edge is the one closest to ${v}_{\tt{start}}$, (2) then append start and goal node as ``$/{v}_{\tt{start}}\hskip 2.84526pt{v}_{\tt{goal}}="$ and (3) then append the full path repeating start and goal node, "${v}_{\tt{start}}\hskip 2.84526ptv_{i_{1}}\dots v_{i_{l-1}}\hskip 2.84526pt{v}_{\tt{goal}}"$. Note that (1) and (2) make up the prefix $\bm{p}$, which the model does not learn to predict. Then, (3) is the target sequence that the model aims to learn. The vocabulary size is thus given by $N+3$, where we add entries for the special tokens ``|"conditional``"``|"` ` | ", ``/"``"``/"` ` / " and ``="``"``="` ` = ". When using the pre-trained models GPT2 we use the tokenizer that was employed for pre-training, in this case the Byte-Pair tokenizer (Radford et al., 2019).
@@ G.2 Models
When training Transformer models from scratch, we use a small model consisting of $n_{\text{layers}}=12$ blocks with embedding dimension $e_{\text{dim}}=384$, $n_{\text{heads}}=6$ attention heads and MLP expansion factor $e=4$, coined GPT-Mini. For pre-trained models, we consider GPT2-Large with $n_{\text{layers}}=36$, $e_{\text{dim}}=1280$, $n_{\text{heads}}=20$ and expansion factor $e=4$ (Radford et al., 2019). To further evaluate purely recurrent models, we perform experiments with the recent Mamba model (Gu & Dao, 2023). We train the Mamba models from scratch with $12$ layers and embedding dimension $784$. We train all the models with the AdamW optimizer (Loshchilov & Hutter, 2019). For models trained from scratch we use a learning rate of $\eta=0.0005$ while for pre-trained models we use a smaller one of $\eta=0.0001$. In both cases we use weight decay of strength $0.01$. Models from scratch are trained for up to $500$ epochs in order to ensure convergence. Pre-trained models require less training time and we usually fit the training data perfectly after $10$ epochs.

### v3

@@ F.2 Clever Hans Cheating Accuracies
GPT-Mini 99.799.799.7 100100100 100100100 99.899.899.8 0.00.00.0
GPT2-Large 99.899.899.8 99.799.799.7 100100100 99.899.899.8 0.00.00.0
Mamba 97.697.697.6 98.398.398.3 99.599.599.5 95.995.995.9 0.00.00.0
@@ F.3 More Detailed Accuracies
GPT-Mini 49.849.849.8 49.149.149.1 19.119.119.1 8.18.18.1 0.00.00.0
GPT2-Large 48.948.948.9 49.249.249.2 19.419.419.4 10.310.310.3 3.53.53.5
Mamba 48.548.548.5 48.748.748.7 20.220.220.2 9.39.39.3 0.00.00.0
@@
GPT-Mini 99.799.799.7 99.899.899.8 100100100 99.899.899.8 0.00.00.0
GPT2-Large 99.999.999.9 99.999.999.9 99.699.699.6 99.899.899.8 99.999.999.9
Mamba 98.598.598.5 96.296.296.2 99.199.199.1 99.599.599.5 0.00.00.0
@@ F.4 Arithmetic Tasks
while keeping all other aspects of the training pipeline (such as the targets and the objective) the same as in standard training. We again train a GPT-Mini-style transformer with AdamW using a learning rate of 555e-444. We display the resulting test accuracies for both standard, reversed and teacherless training in Fig. 6, as a function of training iterations. Here we follow previous works and at every iteration, sample with replacement from all the possible 333-digit configurations (aside from a test set put aside before). We can indeed see that teacherless training leads to more sample-efficient learning compared to the standard format, but to slightly less efficient learning compared to the reverse format. This again hints at the fact that some form of CleverHans cheating is picked up when learning addition.
In this case though, it is more difficult to precisely characterize the form of the cheat in contrast to the path-finding problem. One possible form of cheating may be that in addition, knowing the leading digits of the answer can provide information about the subsequent digits. For example, when we add two single-digit numbers (picked uniformly at random), knowing that the $10^{\prime}th$ place is a 111 rather than a 0, tells us that the 0th place is more likely to be a smaller digit like 0.
@@ G.1 Tokenization
[…]
@@ G.2 Models
When training Transformer models from scratch, we use a small model consisting of $n_{\text{layers}}=12$ blocks with embedding dimension $e_{\text{dim}}=384$, $n_{\text{heads}}=6$ attention heads and MLP expansion factor $e=4$, coined GPT-Mini. For pre-trained models, we consider GPT2-Large with $n_{\text{layers}}=36$, $e_{\text{dim}}=1280$, $n_{\text{heads}}=20$ and expansion factor $e=4$ (Radford et al., 2019). To further evaluate purely recurrent models, we perform experiments with the recent Mamba model (Gu & Dao, 2023). We train the Mamba models from scratch with 121212 layers and embedding dimension 784784784. We train all the models with the AdamW optimizer (Loshchilov & Hutter, 2019). For models trained from scratch we use a learning rate of $\eta=0.0005$ while for pre-trained models we use a smaller one of $\eta=0.0001$. In both cases we use weight decay of strength 0.010.010.01. Models from scratch are trained for up to 500500500 epochs in order to ensure convergence. Pre-trained models require less training time and we usually fit the training data perfectly after 101010 epochs.